# Language Model Council: Benchmarking Foundation Models on Highly Subjective Tasks by Consensus

### v1

@@ 4 Results and Findings
council (by majority vote) 73.7% 75.3%
council (by mean pooling) 74.7% 68.5%
council (no aggregation) 90.5% 52.3%
@@
Table 1: The LMC promotes equal participation as respondents and judges. The Council EI rank and scores are derived from the “council (no aggregation) setting,” where ratings from all LLMs are tallied verbatim, without aggregation or modification. Under various aggregation algorithms, the council is more separable and more consistent than individual LLM judges.
4.1 On Respondents
Table 1 shows the main results of our Council-based EI study. In stark contrast to other leaderboards, Qwen-1.5-110B (#20 on Chatbot Arena) attains the highest ELO score on our EI task, followed by GPT-4o (#1 on Chatbot Arena) in second place. This may be due to the use of its predecessor, Qwen-1.5-32B, as the only common reference model, giving Qwen-1.5-110B a family-specific advantage. A qualitative analysis in Section 5.2 investigates this further. Within other model families, larger models generally rank higher than their smaller, older counterparts, with Gemini-1.0-pro (ranked 4th) versus Gemini-1.5-pro (ranked last) being a notable exception.
Losers ignore word limits. Despite a suggested 250-word limit, a handful of models consistently generate much shorter responses even with ample max_new_tokens. The top 14 models used most of the allowed word length (220+ words), while all models in the bottom 4 averaged less than 200. Notably, Gemini-1.5-pro, despite being the successor to Gemini-1.0-pro, places last with responses averaging 115 words, less than half of what was allowed. Adherence to suggested word limits may be an area of improvement for LLMs and something to consider for future length-conscious benchmarks.
4.2 On Judges
Models with lower invariability during calibration (Appendix B) were also less consistent in the main experiment, indicating that calibration can be used to anticipate a judge’s performance.
Variable Correlation
Council EI Score -0.04
Consistency 0.74
Conviction -0.44
Contrarianism -0.78
Polarization 0.86
@@ 4.2 On Judges
Table 2: Spearman correlation between EI score and key judging qualities to separability.
Consistency strongly correlates with separability (Table 2), likely due to reduced noise in ratings resulting in tighter confidence intervals. Polarization also shows strong correlation, indicating respondents are evenly clustered with moderate outliers. qwen1.5-72B-Chat has the highest contrarianism score but only 38.9% separability. Notably, claude-3-opus achieves the 2nd highest separability despite expressing strong preferences only 0.1% of the time. Conversely, llama-3-8b-chat has the highest conviction, but only 31.1% separability. These findings suggest that the quality of votes is more critical for separability than the strength or incongruity of opinion. Finally, the Council’s EI score is weakly correlated with key judging qualities, indicating that effective judging and task completion are distinct skills.
Table 12 shows that 12 out of 20 LLMs show positive self-enhancement bias, with llama-3-* models exhibiting the highest levels (+0.11 for 70b, +0.21 for 8b). Interestingly, six out of 20 LLMs exhibit negative self-bias. The overall ranking remains similar when self-graded battles are excluded, confirming the effectiveness of using an ensemble of LLMs to mitigate self-bias. While length bias is high, excluding the bottom four models, which were significantly under-length, makes length bias insignificant (Table 3), suggesting that much shorter responses are systematically not preferred.
Models with $>$200 words All models
All votes Consistent votes All votes Consistent votes
Average judge 0.158 0.143 0.502 0.319
Council (majority vote) 0.116 0.129 0.365 0.354
Council (mean pooling) 0.112 0.139 0.592 0.389
Council (no aggregation) 0.125 0.106 0.545 0.347
4.3 Comparison with Human Performance
Human GPT-4o C-A C-M)
@@ 4.3 Comparison with Human Performance
Figure 2: LLM rankings from different systems.
To validate the Council’s evaluations, we conduct a human study mirroring that with the Council: human raters are asked to choose the best of two responses to a given dilemma. We also invite participants to justify their choice based on perceived EI and empathy (adapted from the PETS questionnaire [34]), actionableness, clarity and conciseness. In addition, they have the option to provide a free-text explanation for their choice.
We select nine LLM council members from our pool of 20 for this study, ensuring variety in terms of model size, open- versus closed-source, and company origin. They are claude-3-opus, Qwen1.5-110B, gpt-4o, Llama-3-70b, claude-3-haiku, Llama-3-8b, Qwen1.5-32B-Chat, and gpt-4-0613.
We recruit participants via crowdsourcing on Prolific [^3] https://www.prolific.com/. We randomly sample 120 dilemmas-response triples from the LLM generation. We recruited a total of 102 participants. Each dilemma pair and response were rated by 11 participants on average, with a total of 1343 ratings. Detailed information on our recruitment process, quality control, and participant demographics is provided in Appendix D.
We follow the same methodology described in Section 3.3. As shown in Figure 2, the relative ranking of the systems is consistent between the humans and the LLM council, with small changes. The three top-performing models, as well as the bottom ones, emphasizing a consensus on the ranking of the best and worse models. Compared to the human agreement of 65% on MT-Bench [^4] Agreement between two judges is defined as the probability of randomly selected individuals (but not identical) of each type agreeing on a randomly selected battle., Figure 4 confirms that our EI task is more subjective, with a lower human agreement of 51.9%. Still, the level of agreement between humans and the Council with and without aggregation is roughly the same as the agreement between humans.
4.4 Comparison to Other Leaderboards
Figure 3: Spearman correlation analysis on evaluation benchmarks for 9 LLMs.
Since our case study aims to rank models on a narrow but subjective EI task, high correlation with Chatbot Arena, which measures general capability, is neither intended nor a desired outcome of our project. Instead, Figure 3 shows that the Council successfully exhibits high correlation with the human study, despite the highly subjective nature of our EI task.
@@ 5 Discussion
5.1 Can Judging Consistency be Improved with Oligarchical Councils?
Table 1 shows a wide variety of judging qualities within the council, especially in consistency and separability. In the two-game setup, we collect ratings for models in both positions of pairwise comparisons, which allows us to identify and potentially remove inconsistent ratings before ELO scoring. To explore the dynamics of voting aggregation and the exclusion of inconsistent votes, we formed three oligarchical councils with subsets of LLMs: Flagships, the largest LLM from each organization; Smalls: the smallest LLMs from each organization; and Top-4, the top 4 LLMs according to Chatbot Arena. See Table 9 for a detailed list. The ratings from the main experiment are reused to assess what “would have been” if the sub-council was used instead of the full council.
@@ 5.1 Can Judging Consistency be Improved with Oligarchical Councils?
Figure 4: Separability scores achieved by different council compositions and aggregation methods.
Across all council compositions, the separability of majority aggregation is always higher than the average judge and often higher than the best judge. The council without aggregation achieves the highest separability, probably due to more ratings, which generally means tighter confidence intervals.
However, the smallest council, even with more votes, cannot reach the separability level of the best council with just four judges. Moreover, using only consistent votes means that sometimes  50% of ratings are filtered out, yet separability scores increase.
The council of the smallest models has the highest rate of inconsistent votes and the largest boost in separability when these votes are removed. In all configurations, the top 2 ranked models are always Qwen1.5-110B and GPT-4o, respectively.
To achieve the highest separability and sharpest confidence intervals, especially in low data scenarios, we recommend using no aggregation in LLM evaluation ensembles.
all 0.92 (+0.01) 0.05 (+0.04) 1.0 (+0.48) 0.81 (+0.27) 0.35 (-0.19)
top-4 0.86 (+0.07) 0.04 (+0.03) 1.0 (+0.48) 0.84 (+0.24) 0.25 (-0.13)
Table 5: Changes to key judging qualities when only using consistent votes in the (no aggregation) council.
5.2 Qualitative Analysis: What Makes a Response Preferred Over Another?
Several arena-based benchmarks (ours included) have demonstrated that a clear ranking among LLMs can be established, but there is a dearth of analysis and understanding as to why the rankings are the way they are. For example, platforms like Chatbot Arena do not clarify how factors like feel and style are weighed against correctness [39], and CoT explanations from MT-Bench remain unanalyzed.
We aim to better understand the qualitative aspects of what makes a response to an emotional interpersonal conflict more desirable to better inform how to improve future models. First, we manually examine a random sample of 50 explanations, identifying 38 different reasons for preferences (e.g., “more practical”). The full list is in Appendix E. Next, we used a strong LLM (GPT-4o) to map a larger sample of 1K explanations to these predefined reasons (prompt in Figure 26). The 1K sample includes ratings from all 20 LLM judges. Detailed reason citation frequencies are listed in Figure 19.
We find that LLM judge-provided ratings are almost always based on multiple indicators (4.5 ± 2.4 on average). "More actionable" is the most cited reason, which aligns with the action-oriented framing of our emotional intelligence test. "Structure," "clarity," and "specificity" dominate the top 10 reasons. "More gentle" and "more soft" are cited least, contrasting with "more practical" (#11) and "more authentic" (#12). Longer responses ("more comprehensive” #2, "more detailed" #3) are preferred over brevity ("less verbose" #9).
We also examine feedback from the human study: we find that users generally find that the best responses display emotional intelligence (60.9%), are actionable (55.1%) and clear (52.9%). In contrast, participants reported the best response is concise only 15.9% of the time, suggesting length is less of a determining factor for humans. Moreover, we find little support for empathy: the participants did not find any of the statements in the PETS questionnaire to ring any truer for the winning response. Participants who provided verbal feedback emphasized specificity to the situation, clear examples of how to proceed, and a tone that was not too formal. Full details in Appendix D.
Overall, feedback from human participants and automatically extracted reasons from LLM judge explanations share consistent themes: longer responses that are clear, detailed, and actionable are better when responding to emotional interpersonal conflicts.
Reason 1 Reason 2 Correlation
less verbose more succinct 0.650
better structured more structured 0.584
easier to follow better structured 0.520
easier to follow more structured 0.468
less verbose more direct 0.450
more understanding more empathetic 0.418
more clear better structured 0.387
more direct more succinct 0.349
easier to follow more clear 0.348
more gentle more soft 0.337
Table 6: Top 10 positive correlations.
Reason 1 Reason 2 Correlation
more comprehensive less verbose -0.276
less verbose more detailed -0.227
more comprehensive more direct -0.202
more comprehensive more succinct -0.197
more detailed more succinct -0.196
more comprehensive more focused -0.161
more detailed more direct -0.148
more suggestions … less verbose -0.144
more understanding more actionable -0.139
less verbose more nuanced -0.135
Table 7: Top 10 negative correlations.
5.3 Jury Ablation: What is the Value of the Incremental Judge?
If there are $n$ test prompts and $m$ council members, the cost of evaluating a single LLM is $O(nm)$, or $O(nm^{2})$ for $m$ LLM judges. Unsurprisingly, full participation is costly – as $m$ increases, evaluation costs rise significantly compared to using a single LLM judge.
Councils with full participation achieve the highest separability scores. However, strong oligarchical councils also show competitive separability, suggesting that full participation may not be necessary for sufficiently good rankings. Conversely, councils of smaller models exhibit worse judging quality, emphasizing the need for thoughtful council composition.
To explore the trade-off between efficiency and inclusivity in determining council size and composition, we use Monte Carlo methods to simulate councils of different sizes and compositions, assessing what the rankings would be with different subsets of LLM juries.
We quantify stability by measuring the Mean Variance of the Rank (MVR), defined as the variance of a respondent’s rank averaged over all respondents and random trials. An MVR of 1 means a respondent’s rank is expected to change by 1 position in a new trial. Lower MVR indicates more stable rankings, with MVR of 0 signifying perfect stability. To assess robustness, we observe the impact of adding adversarial judges who vote randomly. Results for 100 random samples for increasing council sizes are shown in Figure 5.
Figure 5: Monte Carlo jury ablation measuring mean variance of the rank (MVR) (left) and separability (right) using majority-aggregated juries. The x-axis represents the number of non-adversarial judges in each jury. Each data point represents the score over 100 randomly sampled jury compositions. The 95% confidence intervals are shown for MVR. For separability, random trials bootstrap confidence intervals directly.
Larger juries have lower MVR, both with and without adversarial judges. Adversarial judges negatively impact both MVR and separability. Larger councils demonstrate consistent robustness to adversarial judges in both MVR and separability, though with diminishing marginal returns, flattening around councils of size 12.
6 Limitations
We only studied single-turn arguments. Our study evaluates EI based on exposure to single, self-contained interactions. In many contexts, EI is effectively applied through iterative discussions. A more interactive setup with dynamic exchanges could create stronger tests for EI.
Cultural and linguistic context. Our study focuses on English articles and English speakers. Further research would be needed to determine the broader applicability of our results to other languages.
Resource allocation can be significant for large councils. While the assumption of equal weight for all council opinions guides this work, effective subsets and variations in judging proficiency highlight complexities. Balancing example quantity and council size requires careful consideration.
Reproducibility. Closed-weight models like GPT-4 pose a reproducibility challenge due to undisclosed updates, causing responses to vary despite fixed settings. To ensure reproducibility, we provide details like the source and hyperparameters used.
Emotional responses are personal, influenced by experiences and social factors. The effectiveness of LLM’s response is best judged by the person experiencing the emotion. Our quantitative evaluation may not fully capture the essence of a first-person judgment of conversational responses.
@@ 7 Conclusion
In this paper, we propose the LMC as a flexible evaluation framework for ranking LLM agents in a fully democratic fashion. The framework is flexible, easily extendable, and its decentralized design is inherently robust to adversarial participants. An in-depth case study applying the Council to a highly subjective emotional intelligence task with 20 state-of-the-art LLMs shows that Council, in a fully decentralized manner, produces highly separable rankings that correlate significantly more with human-established rankings than other benchmarks. Altogether, the LMC establishes a foundation for a consensus-based LLM evaluation framework for highly subjective tasks.

### v2

@@ 4 Results and Findings
LMC (majority vote) 73.7% 75.3%
LMC (mean pooling) 74.7% 68.5%
LMC (no aggregation) 90.5% 52.3%
@@
Table 1: The LMC promotes equal participation as respondents and judges. The Council EI rank and scores are derived from the “council (no aggregation) setting,” where ratings from all LLMs are tallied equally, without aggregation or modification. Under various aggregation algorithms, the council is more separable and more consistent than individual LLM judges.
Table 1 shows the main results of our Council-based EI study, highlighting the following insights:
Qwen-110B outranks GPT-4o in an unexpected upset.
Like other benchmarks, larger models within the same family tend to outrank their smaller or older versions. However, unlike other benchmarks, Qwen-1.5-110B (#20 on Chatbot Arena) scores highest on our EI task, followed by GPT-4o (#1 on Chatbot Arena) [^10] This was the ranking as of May 2024. The Qwen-1.5 models have been dropped from Chatbot Area since the Qwen-2.5 family of models have been added.. This is surprising, as Qwen-1.5-110B does not typically outperform GPT-4o. One possible reason for this outcome is the use of Qwen-1.5-32B as the reference model (Appendix D). Because all of Qwen-1.5-110B’s responses are compared to the responses of a strictly smaller variant of the same family, this could result in an outsized advantage for Qwen-1.5-110B in the evaluation overall. This raises an interesting implication for Arena-style evaluations, as it’s possible that the choice of reference model may introduce a bias favoring its successors in the same arena.
Judges prioritize actionability, clarity, and structure when expressing preferences.
Using chain-of-thought (CoT) prompting, judges provided detailed reasoning for their preferences. We analyzed 1,000 reasoning traces and identified common themes (Appendix G).
Judges disfavor models that produce responses significantly shorter than the suggested word limit.
Despite a suggested 250-word limit, some models generated much shorter responses, even though decoding parameters allowed for longer output. Models that adhered to the limit, using 220+ words on average, performed better, while models in the bottom four positions averaged less than 200 words. Notably, Gemini-1.5-pro placed last with an average response length of just 115 words, far worse than its predecessor Gemini-1.0-pro (4th place) with 228 words on average. LLM Judges bias towards longer responses, but excluding the outliers (models that went well under the limit) made length bias insignificant (Table 6).
@@
Figure 2: Spearman correlation between EI score and key judging qualities across 20 LLM council members.
LLM success in the EI task does not correlate with its judging ability.
Performance on the EI task had only a weak correlation with any of the key judging qualities (Figure 2), which suggests that the ability to perform well in the task and the ability to judge others’ responses are distinct skills.
Consistent judging, neutral voting patterns, and lower contrarianism lead to higher separability.
Our analysis shows that judges with more consistent votes and neutral voting patterns tend to achieve higher separability. Judges with high conviction — those that express strong preferences (A>>B or B>>A) frequently — are negatively correlated with separability, suggesting that overly strong votes tend to introduce noise. For example, claude-3-opus, despite expressing strong preferences only 0.1% of the time, achieves the second-highest separability (72.6%). Higher contrarianism, which measures how often a judge disagrees with the majority decision, also correlates with lower separability. Qwen1.5-72B, the most contrarian judge, has only 38.9% separability.
LLMs are lightly self-biased, but it is mitigated in the full Council.
Out of the 20 LLMs in our study, 12 exhibited positive self-enhancement bias, meaning they rated themselves more favorably than the Council’s overall score for them. When self-graded battles are excluded, the overall rankings remain similar (Figure 9). This indicates that while individual models carry self-enhancement bias, the ensembling of the LMC effectively neutralizes these biases.
Agreement among LMC members, humans, and between the LMC and humans is similar.
Figure 3 shows that the rankings produced by LMC members and humans are consistent, with small variations. Both groups agree on the top-performing models and the lowest-ranked ones. The level of agreement between the LMC and humans is roughly the same as the agreement within human evaluators (51.9%).
Human GPT-4o LMC-A LMC-M
@@
Figure 3: LLM rankings from different benchmarks.
Figure 4: Spearman correlations for different evaluation benchmark scores for 9 LLMs.
The LMC’s rankings align more closely with human evaluations than other benchmarks or individual judges.
The LMC achieves the highest correlation with human-established rankings, outperforming other benchmarks, including those from similar domains like the EI-specific subset of Chatbot Arena (Appendix F) and EQ-Bench Paech (2023) (Figure 4). The LMC outperforms most individual LLM judges in terms of correlation with human rankings. While a few individual LLM judges, like dbrx-instruct, achieve similar correlation scores (e.g., 0.917), they do so with significantly lower separability (50.5% vs. 90.5% for the LMC). This suggests that the LMC’s collective judgment not only aligns more closely with human preferences but also provides clearer distinctions between model performance, making it a more reliable method for ranking models in our case study.
Additional findings.
The 20x20 LLM interactions generate a wealth of data, some of which is not included in the main paper for brevity. Additional insights can be found in Appendix B.
@@ 5 Discussion
One of the most important questions about using a LMC is whether it is worth the trouble. Collecting more opinions costs more energy, and parsing, tallying, and storing them add logistical costs. What is the value of another opinion, much less everyone’s opinion?
From a cost perspective, a closely related question is how many examples we should collect judgements for to begin with. Using a smaller test set simplifies stakes for human reviewers — it is far easier to read and remember 10 examples than 10,000 – while also reducing costs.
If we assume that the main goal is to establish good relative rankings, we narrow the focus of quantifying the success of a council by measuring the significance and stability of the final ranking. For significance, we look at separability (as in the main experiment), which measures how well models can be distinguished based on non-overlapping confidence intervals. For stability, we create a new metric, Mean Expected Rank Variance (MERV), defined as the expected ordinal swing of the average respondent’s rank (Appendix H). A MERV of 3 means an average respondent’s rank is expected to change up to 3 positions in a new trial. Lower MERV indicates that the benchmark has more stable rankings, with MERV of 0 signifying perfect deterministic-like stability.
5.1 Monte Carlo simulations
To study the dynamics of separability and stability with differently-sized councils and test sets irrespective of the inclusion of any specific LLM, we use a Monte Carlo procedure to simulate many random hypothetical councils and test sets.
Our Monte Carlo simulation procedure is as follows: 1) For a given council size $c$ and test set size $t$, randomly sample $c$ LLMs to form a council $C$ and $t$ examples to form a test set $T$. Sampling is performed with replacement. 2) Find the associated judgements from the main experiment for the specific ($C$, $T$) configuration to determine the scores and relative rankings for all LLMs. 3) Repeat for 100 trials [^11] We use 100 to be consistent with the number of rounds of bootstrapping in the main experiment. For calculating separability in Monte Carlo simulations, the trials themselves are used to bootstrap confidence intervals directly.. 4) After all 100 trials are complete, tally the results: for stability, observe fluctuations in rankings for each LLM to compute MERV, and for significance, report the mean separability. Figure 5 shows results for a sweep $c\in\{1,3,5,\dots,19\},\quad t\in\{10,20,30,\dots,100\}$.
(a) MERV
(b) MERV gradients
(c) Separability $\mu$
(d) Separability $\mu$ gradients
Figure 5: Measurements of rank stability (MERV) ((a) and (b)) and separability ((c) and (d)) averaged over 100 randomized trials for various numbers of judges and examples. (a) and (c) display raw metric values while (b) and (d) display the gradient magnitude (colors) and direction (arrows). The gradient calculation follows a Manhattan distance approach where row-wise and column-wise gradients are linearly combined to reflect the discreteness of changes between adjacent squares, highlighting the incremental impact of adding another judge or more examples.
A nuanced trade-off between the size of the test set and the number of judges.
Both separability and stability improve as the number of test examples and the number of LLM judges increase, with the best scores achieved when both are maximized. However, based on the gradient maps for MERV (Figure 5(b)) and separability (Figure 5(d)), the added benefit of including either an additional judge or more test examples diminishes significantly in a concentric shape that starts  50 examples and  9 judges. The gradients in this darker zone indicates where the utility of any additional opinion—whether in the form of new test data or another LLM judge—becomes marginal.
Figure 6: MERV and separability for Monte Carlo simulations for ($t$=30) with adversarial judges.
Larger councils are more robust to adversarial judges, with diminishing marginal utility.
As the size of the LMC grows to include many different LLMs, it may become difficult to verify the quality of every member, particularly on subjective tasks. In our Monte Carlo procedure, we experiment with a simulation setting with adversarial judges. An adversarial judge is a fake LLM council member that returns ratings at random [^12] There are other adversarial algorithms that can be used such as always voting for the first position that we did not explore.. In Figure 6, we find that on both separability and stability, larger councils reduce the negative impact of adversarial LLM judges. We find that this robustness continues to increase as the size of the council grows, even when maintaining the same ratio of adversarial judges to real judges, albeit with diminishing marginal utility.
What is the value of the incremental judge?
With respect to stability and separability, it depends. For small test sets, the value of additional test examples outweighs the benefit of more judges, but when test data is abundant test data, the value of an additional judge is higher. If there are more than 20-30 examples in the test set, adding an additional judge provides greater value, but for less, it is better to get opinions on an additional batch of examples with the current set of judge(s).
Overall, larger councils are more robust to adversarial judges so the bigger your council is, the less concerned you need to be about a particularly bad judge from wreaking havoc on your study, and you don’t have to be as picky about your council members.
5.2 Oligarchical councils
If a task does benefit from multiple perspectives and the evaluation budget allows for a limited number of opinions, whose opinions should we include? Verga et al. (2024) suggests replacing a single large judge with 3 smaller judges. [^13] Council composition is likely a task-specific optimization; On HotpotQA, an ensemble of LLM judges was worse than using Claude Haiku alone Verga et al. (2024). We explore a similar but reverse question: can a subset of judges (or a single judge) effectively represent the fully democratic LMC?
We compare the rankings of three hand-curated sub-councils (Table 4): flagships, smalls, and top-4. We refer to the sub-council as an “oligarchical” council since we rely on a small subset of participating LLM judges to determine the ranking of every LLM, akin to oligarchies in human society.
Our analysis finds that full council participation is neither necessary to achieve rankings that align with human rankings nor maintain good separability. Smalls, a council composed of the smallest LLMs, achieves a respectable separability of 71%, which is already better than the average judge (53.3%). It also achieves a Spearman correlation of 0.88 with human rankings, only slightly lower than the full council’s correlation of 0.92 (Figure 23).
Still, it’s worthwhile to be thoughtful about council composition. For instance, top-4 achieves the same correlation with human evaluations as smallest (0.88), but with notably higher separability (79%). While smaller councils consisting of smaller LLMs can perform reasonably well, there are other factors to consider like separability, robustness, human alignment, and bias mitigation.
@@ 5.2 Oligarchical councils
Figure 7: Separability scores achieved by different council compositions and aggregation methods. Higher separability is better.
@@ 6 Conclusion
In this paper, we introduce the Language Model Council (LMC), a flexible, decentralized evaluation framework for ranking LLM agents through democratic participation. Applying the LMC to an emotional intelligence task with 20 LLMs, we demonstrate that the LMC can produce highly separable rankings that align more closely with human judgments than other benchmarks or individual judges. Through both Monte Carlo simulations and hand-curated sub-councils, we find that while larger councils provide benefits, they are incrementally diminishing, and the majority of key qualities—such as ranking significance, stability, and alignment with human evaluations—can be achieved with a smaller, well-chosen ensemble of judges. As humans increasingly rely on LLMs to evaluate other LLMs, we hope the LMC framework, along with insights from our case study, offers a valuable foundation for developing reliable LLM evaluations, even for highly subjective tasks.
Limitations
Generalizability of the LMC.
Although we present one detailed case study focused on EI, the Language Model Council (LMC) framework is broadly applicable to a wide range of open-ended tasks. The core mechanism of tallying preferences through arena-style pairwise comparisons is inherently adaptable to various types of prompts and tasks Chiang et al. (2024). However, framing new subjective tasks, such as those related to aesthetics or politics, in a form suitable for technical evaluation still requires careful design. In our case study, we were responsible for the technical formulation of the EI task, and the task examples were seeded from EmoBench Sabour et al. (2024), a human-crafted dataset that we chose.
The least generalizable aspect of the LMC is likely the first step: formulating the test set in a collaborative way. For tasks with fixed or human-authored test sets, it may be undesirable or unclear how to implement participation from multiple LLMs. In such cases, this step could be omitted or delegated to a single strong LLM. Whether any LLM can generate meaningful test sets for narrow tasks in a fully unsupervised, domain-generic manner—subjective or otherwise—remains an open area of exploration Lhoest (2024).
Single-turn interactions and English-only evaluation.
Our case study evaluates EI based on single, self-contained interactions and was conducted entirely in English. However, many tasks may be better assessed through extended conversations, multiple sessions, multiple modalities, or in multiple languages, to reflect a broader range of human interaction dynamics, all of which are not explored in this paper.
Reproducibility challenges.
LLMs are inherently stochastic, meaning the same model can produce different ratings even with temperature set to 0 Chann (2023). Reproducibility is further complicated by closed-weight models like GPT-4, which may receive undisclosed updates. All responses in our study were collected in May 2024, but serverless providers, like Together, may update their APIs or discontinue support for certain models, as happened with Qwen-1.5-32B (replaced by Qwen-2.5). For open source models, changes in deployment hardware or GPU configuration can introduce additional variability, making exact replication of results difficult.
LLM statelessness.
We assume that LLMs, being memory-less, can serve as both respondents and judges simultaneously. In contrast, humans would struggle to judge their own responses impartially due to memory retention. As LLMs evolve to incorporate memory—such as retaining recent prompts and responses OpenAI (2024c)—the risk of self-enhancement bias may increase. If LLMs begin to remember their own responses during evaluation, disabling self-grading may become the default approach to ensure fairness.
Inclusive democracy does not guarantee fairness.
While the LMC mitigates self-enhancement and position bias, it remains vulnerable to systemic biases within the evaluation framework itself. For instance, we speculate that using a single reference model may introduce an advantage for successors of the reference model in the arena. This specific aspect could be mitigated by using multiple reference models, but it can be challenging for new test data to compensate for a bias in fundamental evaluation procedures.
Diversity of opinions in LLMs versus humans.
The question of whether the distribution of an ensemble of LLM judgments is a good approximation for general human diversity Dong et al. (2024); Hosking et al. (2023) is secondary to our focus on the alignment of final rankings, which is the aggregated expression of the collection of abundantly dissenting opinions (from humans or LLMs), and thus where we assert the utility of the LMC’s methodology. In our EI case study, we found that the LMC’s final ranking aligned better with human preferences compared to other benchmarks and the rankings produced by nearly all individual LLM judges.
We also recognize that the human judgments collected in our study may not fully represent the diversity of opinions within the broader population Elangovan et al. (2024), nor do they reflect authentic, first-hand judgments. Emotional responses are highly individual, and real interpersonal conflicts are shaped by personal experiences and social factors that may be difficult for anyone other than the person experiencing the conflict to fully evaluate. For both humans and LLMs, we can only make a deliberate effort to gather judgments from some accessible variety of relevant profiles. This approach mirrors the rationale behind the design of the LMC in our case study, which was similarly formed with variety in mind.